4 Zn + 10 HNO3     -->    4 Zn(NO3)2  + N2O   + 5 H2O


How many g of HNO3 would be needed to produce 34.6 L N2O?

How many mol N2O would be produced from 134.5 g Zn reacting with excess HNO3?If 125 g H2O were produced, how many g Zn were used?

How many g Zn would be needed to produce 2.55 x 1024 molecules N2O?





1
Expert's answer
2021-04-16T04:59:03-0400

1)

Assuming STP conditions,

"m(HNO_3)=34.6L(N_2O)\times\frac{1mol(N_2O)}{22.4L(N_2O)}\times\frac{10mol(HNO_3)}{1mol(N_2O)}\times\frac{63.01g(HNO_3)}{1mol(HNO_3)}=973g"

Answer: 973 g


2)

"n(N_2O)=134.5g(Zn)\times\frac{1mol(Zn)}{65.38g(Zn)}\times\frac{1mol(N_2O)}{4mol(Zn)}=0.5143mol"

Answer: 0.5143 mol


3)

"m(Zn)=125g(H_2O)\times\frac{1mol(H_2O)}{18.015g(H_2O)}\times\frac{4mol(Zn)}{5mol(H_2O)}\times\frac{65.38g(Zn)}{1mol(Zn)}=363g"

Answer: 363 g


4)

"m(Zn)=2.55\cdot10^{24}(molecules\;N_2O)\times\frac{1mol(N_2O)}{6.022\cdot10^{23}(molecules\;N_2O)}\times\frac{4mol(Zn)}{1mol(N_2O)}\times\frac{65.38g(Zn)}{1mol(Zn)}=1110g"

Answer: 1110 g